Find the equations of the tangent and the normal of y = 3x2-2x+1 at point (1,2).


1
Expert's answer
2022-03-04T09:29:23-0500
"y'=(3x^2-2x+1)'=6x-2"

Point (1,2)


"slope_1=m_1=y'(1)=6(1)-2=4"

"slope_2=m_2=-\dfrac{1}{y'(1)}=-\dfrac{1}{4}"

The equation of the tangent line is


"y=4(x-1)+2"

"y=4x-2"

The equation of the normal line is


"y=-\dfrac{1}{4}(x-1)+2"

"y=-\dfrac{1}{4}x+\dfrac{9}{4}"